# Reconstructing Movement from Sparse Samples: Enhanced Spatio-Temporal Matching Strategies for Low-Frequency Data

### 10 Mar 2026 v1

@@ Abstract
This paper explores potential improvements to the Spatial-Temporal Matching algorithm for matching the GPS trajectories to road networks. While this algorithm is effective, it presents some limitations in computational efficiency and the accuracy of the results, especially in dense environments with relatively high sampling intervals. To address this, the paper proposes four modifications to the original algorithm: a dynamic buffer, an adaptive observation probability, a redesigned temporal scoring function, and a behavioral analysis to account for the historical mobility patterns. The enhancements are assessed using real-world data from the urban area of Milan, and through newly defined evaluation metrics to be applied in the absence of ground truth. The results of the experiment show significant improvements in performance efficiency and path quality across various metrics.
@@ 1 Introduction
The widespread use and availability of Global Positioning System (GPS) technologies now enables the continuous monitoring of vehicles or people across both time and space. However, due to intrinsic errors and inaccuracies in positioning systems, raw trajectory data often contain noise, missing segments, and spatial misalignment. Several factors may negatively affect the accuracy of the individual GPS points recorded, such as signal blockage in dense urban environments and inconsistency of GPS receivers. As a consequence, the recorded trajectory may inevitably deviate from the true path taken on the road network [5]. In addition, by nature, GPS data provide only a partial view of real-world movement: continuous motion is observed only at discrete time intervals, capturing therefore just a subset of the true trajectory. This limitation becomes more pronounced when the data are sampled at low frequency, resulting in large temporal gaps between consecutive observations. In such cases, important details of the trajectory may be lost, making it difficult to precisely reconstruct the true travel path [10].
To address these limitations, a preprocessing stage is required to facilitate better interpretation of the raw location data, with one of the most essential steps being Map Matching [2]. The Map Matching technique is utilized to match the potentially noisy GPS observations of spatial trajectories with their real position on the underlying transportation network, identifying the correct road segment on which a moving vehicle or user was traveling, and their position on that segment [9]. It acts as a fundamental step in various location-based applications such as navigation, tracking of vehicles, traffic monitoring, and intelligent transport systems [2, 16].
@@
In this context, one widely cited method for low-frequency GPS data is the ST-Matching, proposed by Lou et al. in 2009 [6]. This algorithm is a weighted graph technique, i.e., a specific case of state transition models [2]. Indeed, it forms the matched path based on the weights of a candidate graph, whose nodes represents the point locations on the road map that are candidates for being the true position of the GPS record, and whose edges connect pairs of consecutive points and are weighted based on spatial proximity and a temporal function. As a result, a Directed Acyclic Graph (DAG) is formed, and the target becomes finding the longest path on this weighted graph, i.e., the sequence of candidates that maximizes the total weight utility. Despite its effectiveness, the method struggles with complex road topologies, noisy GPS data, and scalability issues when processing large datasets.
The goal of our work is to investigate how the ST-Matching algorithm can be improved both in terms of its performance efficiency and the quality of its results. To address this, a set of modified versions of the method is proposed, each designed to overcome specific limitations of the original implementation and enhance the overall performance. In particular, the ST-modified Matching Map Matching is introduced, refining the original method by improving the candidate point selection method, the spatial analysis component, and the temporal component to better reflect realistic movements, travel time, and speed. Moving on to the realm of low-frequency data, we investigated the potential of adding a behavioral score based on historical data to the algorithm, leading to the design of a new behavioral score and the development of the STB-Matching method.
In addition, this paper presents an evaluation framework designed to assess the performance of the modified algorithms. Several metrics are proposed to examine the methods from different perspectives, including computational efficiency, matching accuracy, adherence to topological constraints, and execution speed. Throughout them, a comprehensive comparative analysis is performed to assess the improvements of the new methodologies with respect to the baseline method. The performance and results of this algorithm are also compared to the baseline on low-frequency data. The proposed approach is validated through a case study in Milan, a key metropolitan area in Northern Italy. The results show clear improvements over the baseline ST-Matching implementation and highlight which modifications are most effective under various operating conditions.
@@ 2.1 Prior Definitions
Considering the the road network graph $G(V,E)$, the path $P_{ij}$ between two vertices $V_{i}$ and $V_{j}$ is a set of connected road segments $e_{1}\rightarrow e_{2}\rightarrow\dotsi\rightarrow e_{k}$, where $V_{i}$ is the starting point of the first edge $e_{1}$, and $V_{j}$ is the ending point of the last edge $e_{k}$.
@@ 2.2 Overview of the Baseline Method
The observation probability measures the likelihood that GPS sample point $p_{i}$ corresponds to a candidate point $c_{i}^{s}$, given the Euclidean distance between them. This likelihood assumes that GPS error has a normal distribution, typically with the mean of zero and a standard deviation ($\sigma$), which is set to $\sigma=20$ based on the empirical evaluations in the original work [6]. The observation probability is calculated using the probability density function of such a distribution as follows:
@@
After spatial and temporal analysis, for each trajectory $T$, it is possible to create a Trellis candidate graph [11]. The set of its nodes includes the candidate points for each GPS sample $p_{i}$, and the set of its edges is made of links connecting candidates of consecutive points. Formally, the Trellis candidate graph is a Directed Acyclic Graph defined as $G_{T}=(V_{T},E_{T})$ with:
@@ 2.3 Proposed Modifications to ST-Matching
    Dynamic candidate preparation, which affects the search area for the candidate points using a dynamic buffer and through GPS accuracy. It is discussed in Section 2.3.1
@@
    Dynamic observation probability, which modifies the first component of the spatial analysis again based on data accuracy, and is presented in Section 2.3.1
@@
    Redesigned temporal score function, which replaces entirely the temporal function of the original algorithm and is detailed in Section 2.3.2
@@ 2.3.2 Redesigned Temporal Function
$v_{avg}(i)=\frac{\text{dist}_{e}(p_{i},p_{i-1})}{\Delta t_{obs}(i)},$ (8)
$\Delta t_{est}(i,s,t)=\frac{\text{dist}_{n}(c_{i}^{s},c^{t}_{i-1})}{v_{avg}(i)}.$ (9)
@@
The speed penalty is designed to compare the average speed estimated from the trajectory data $v_{\text{avg}}$, with the average speed limit of the road segments in a candidate path $v_{\text{lim}}$, and penalizes only over-speeding when the estimated average speed goes beyond the average speed limit. The average observed speed $v_{\text{avg}}$ is defined as in Eq. 8 and the average speed limit $v_{\text{lim}}$ is defined as follows for each candidate pair $c_{i-1}^{t},c_{i}^{s}$ of consecutive points $p_{i-1},p_{i}$:
@@
where $E_{i,s,t}=E_{(i-1,s)\rightarrow(i,t)}$ is the set of edges of the road network included in the shortest path between candidate $c_{i-1}^{t}$ pf $p_{i-1}$ and candidate $c_{i}^{s}$ of $p_{i}$. Similarly to the travel time penalty function, the speed penalty is defined with an exponential shape, as in Equation 12. It keeps the value of 1 when the average speed is below the limit, and drops as it goes beyond the limit.
@@
Final considerations regard the speed variation penalty. Assuming a moving object on the road network of a city, throughout a trajectory, it passes many road segments and intersections, while prefering to follow relatively direct routes rather than changing the road at every intersection. As a result, the speed limit along a real trajectory should not fluctuate dramatically. Speed variation penalty is introduced to capture this behavior. The penalty is defined based on the ratio of the standard deviation to the mean of speed limits along the candidate path. Considering two candidates, $c_{i-1}^{t}$ and $c_{i}^{s}$, and their candidate path $E_{i,s,t}$, we define the mean and standard deviatiton of the speed limits in the candidate path, and thus the speed variation penalty, as follows:
@@
After the calculation of these three components, the final temporal score function can be computed as the product of these penalties:
@@ 2.3.3 Behavioral Analysis
The computation of the behavioral score is as follows. Using historical GPS trajectores of a specific period of time, we perform the candidate point selection stage of the algorithm using the dynamic buffer around all the GPS points in the historical dataset. Each candidate point is mapped to a road segment of the network graph. For every instance a candidate point is associated with a given edge, a counter for that edge is incremented by one. As a result, each edge in the network will contain a usage score representing how often it has probably been used in the past trajectories. To deal with the wide range and skewness in edge scores, we apply a logarithmic normalization to scale the values between 0 and 1:
@@ 2.3.4 Final Score Functions
As a result of all the analyses and modifications discussed so far, we can define the score functions of two versions of the algorithm. The modified ST-Matching uses a score function combining Equations 3 and 16: the buffer and $\sigma$ parameters used to calculate Equation 3 are the dynamic versions, while the temporal function is modified according to Equation 16.
@@ 2.4.1 Efficiency Metrics
The efficiency metrics assess the computational efficiency of the algorithm. By evaluating these metrics, we can quantify the overhead introduced by different candidate selection strategies (e.g., fixed vs. dynamic buffers) and identify trade-offs between algorithmic complexity and matching accuracy. Specifically, we look at the following quantities per each trajectory $T$ of the dataset:
@@ 2.4.3 Topology Metrics
  • Number of revisited edges: The number of times the same edge in the road network graph is used more than once for the same trajectory. Frequent revisits are typically physically implausible for standard transit and suggest that the algorithm is “trapped” oscillating between candidate edges due to noise.
@@
  • Number of loops: Defined by the number of polygons formed by the matched path. While naturally present in roundabouts, an excess of loops identifies topological instability and physically implausible backtracking caused by high GPS noise or weak temporal constraints.
@@ 2.4.4 Speed-Based Metric
The speed-based metric evaluates how realistic the reconstructed travel speed is compared to the expected speed estimated from the GPS samples. The speed estimated by the GPS data is treated as a reference. The metric is calculated as the speed relative error.
@@ 3.1 GPS Trajectory Dataset
The trajectory observations analyzed in this study are located in the Milan urban area and its surroundings and refer to the period from December 1st to December 20th, 2019. The initial dataset is subjected to a filtering process to guarantee the high quality of the location data and its consistency with the objectives of the study. While the focus of the analysis was on the urban area of Milan, the dataset contains many records outside this area, probably related to daily commuters. Therefore, the first filter is to exclude the records outside the urban area, which is done using the official shapefile of the Milan urban area from the open data portal of the municipality [3]. The sampling interval of the GPS records is quite high in most cases, with the average time interval of 16 seconds over the whole dataset, which means that this is a case of high-frequency GPS data. In order to make sure that the trajectories that are potentially going to be used later in Map Matching are long enough, only the ones with at least 10 GPS records throughout the trajectory are kept. Finally, in this study, the focus is on the vehicular mobility on the drivable road network, which means trajectories corresponding to walking users must preferably be disregarded. This is achieved by calculating an approximate average speed for each trajectory using the direct distance between each two adjacent points and the time interval between them, and keeping the trajectories with an average speed of at least 6km/h to make sure there are no walking users in the dataset. In the end, the final dataset is composed of $10546$ trajectory IDs corresponding to motor vehicles moving in the city of Milan.
@@ 3.2 Road Network Description
The road network used in this paper covers the urban area of Milan, Italy. It is retrieved from OpenStreetMap [8] using the OSMnx Python library [1]. The network type drive is used in this case, which includes only drivable roads. The retrieved graph is directed and multi-edge, meaning multiple edges may exist between a pair of nodes.
@@
The speed attribute of the graph edges showed missing or non-numerical values, a common limitation in the OpenStreetMap database. Since this attribute is essential for the temporal analysis of Map Matching, this issue is addressed in several steps. First, in cases where there are no numerical values, but instead a road type code, it is possible to infer information about the speed limits for the corresponding code from the OpenStreetMap Wiki [7]. In few cases where there is a list of two values for the speed limit, the greater values are kept. Later, dealing with missing values, assuming that different segments of the same street usually have the same speed limit, we use the name attribute (i.e., the street name) to fill in the missing values if other segments of the street with the same name contain a maximum speed value. Finally, for the rest of the missing values, another attribute called highway (i.e., the highway type) is used considering the mode of maximum speed values for each road type. The final distribution of the speed limits of the network edges comprises a majority of road segments of 30 km/h (56.0%), followrd by many road segments of 50km/h (38.1%), and the remaining with other velocities.
@@ 3.3 Results
To simulate a low-frequency trajectory, another subsample of 100 trajectories is considered and limited: for each trajectory from the first GPS point we keep the poits with at least a 2 minute time interval from the previous one. The low-frequency trajectories are then used to compare ST-Matching with its modified version including behavioral analysis. The method evaluation is performed exploiting the proposed evaluation framework, but also comparing the results obtained with the synthetic low-frequency data with their original counterpart, which is used as a ground truth. Moreover, for the behavioral ST-Matching, the edge score of the road network is computed using a different subsample of size 400 of the trajectory dataset.
@@
This section compares the result of Map Matching between the original ST-Matching and the modified version which includes dynamic buffer, dynamic observation probability, and the new temporal function. The comparison is carried out through the metrics mentioned in Section 2.4 based on the distribution of the metric values and the statistical significance expressed by a t-test.
@@
It is observed in Figure 7 that the speed relative error, similar to other metrics, signals an improvement in the results and potentially more realistic paths regarding the temporal aspects of the Map Matching process.
@@
The same comparison across all metrics will be presented in this section using the lower-frequency subsample of the dataset. A part of the dataset is used to calculate the edge score. Figure 10 illustrates the distribution of the edge score on the road network of the urban area of Milan using the available dataset. It can be observed that the roads directed towards the suburban areas and the central rings are generally darker, meaning that they have higher scores and have been used more frequently according to the available dataset.
@@
As expected, efficiency metrics are positively affected by the new algorithm due to the dynamic buffer that results in less candidate points and less runtime. The result of the t-test, as written on the graphs in Figure 11, confirms this effect.
@@
Finally, speed relative error also does not show any significant differences between the two algorithms. It is clearly observed in Figure 14 that the distribution of the values and the t-test results indicate almost no change in this metric.
@@
Finally, we compare the results of the path enhancement using the synthetic low-frequency data and their original counterpart, assuming the matched path of the latter as the ground truth and checking how close are the results of ST-Matching and STB-Matching using the former. In order to do so, the edges of the final matched paths by ST-Matching and STB-Matching are compared with that of the modified ST-Matching to see the number of edges they have in common. As summarized in Table 3, among 100 trajectories, in 72 cases ST-Matching and STB-Matching had the same number of edges in common with the modified ST-Matching results. In 15% cases STB-Matching is better, and in 13% cases ST-Matching is the winner. This reveals that improving the performance of ST-Matching algorithm for low frequency GPS data is a demanding job and might need more investigation.
@@
Table 3: Comparison of overlap between ST-Matching and STB-Matching with the modified ST-Matching results.

### 11 Mar 2026 v2

@@ Abstract
This paper explores potential improvements to the Spatial-Temporal Matching algorithm for aligning the GPS trajectories to road networks. While this algorithm is effective, it presents some limitations in the computational efficiency and the accuracy of the results, especially in dense environments with relatively high sampling intervals. To address this, the paper proposes four modifications to the original algorithm: a dynamic buffer, an adaptive observation probability, a redesigned temporal scoring function, and a behavioral analysis to account for the historical mobility patterns. The enhancements are assessed using real-world data from the urban area of Milan, and through newly defined evaluation metrics to be applied in the absence of ground truth. The results of the experiment show significant improvements in performance efficiency and path quality across various metrics.
@@ 1 Introduction
The widespread use and availability of Global Positioning System (GPS) technologies now enables the continuous monitoring of vehicles and people across both time and space. However, due to intrinsic errors and inaccuracies in positioning systems, raw trajectory data often contain noise, missing segments, and spatial misalignment. Several factors may negatively affect the accuracy of the individual GPS points recorded, such as signal blockage in dense urban environments and inconsistency of GPS receivers. As a consequence, the recorded trajectory may inevitably deviate from the true path taken on the road network [5]. In addition, by nature, GPS data provide only a partial view of real-world movement: continuous motion is observed only at discrete time intervals, capturing therefore just a subset of the true trajectory. This limitation becomes more pronounced when the data are sampled at low frequency, resulting in large temporal gaps between consecutive observations. In such cases, important details of the trajectory may be lost, making it difficult to precisely reconstruct the true travel path [10].
To address these limitations, a preprocessing stage is required to facilitate better interpretation of the raw location data, with one of the most essential steps being Map Matching [2]. The technique is utilized to match the potentially noisy GPS observations of spatial trajectories with their real position on the underlying transportation network, identifying the correct road segment on which a moving vehicle or user travels, and their position on that segment [9]. It acts as a fundamental step in various location-based applications such as navigation, tracking of vehicles, traffic monitoring, and intelligent transport systems [2, 16].
@@
In this context, one widely cited method for low-frequency GPS data is the ST-Matching, proposed by Lou et al. in 2009 [6]. This algorithm is a weighted graph technique, i.e., a specific case of state transition models [2]. Indeed, it forms the matched path based on the weights of a candidate graph, whose nodes represents the point locations on the road map that are candidates for being the true position of the GPS record, and whose edges connect pairs of consecutive points and are weighted based on spatial proximity and a temporal function. As a result, the target becomes finding the longest path on this weighted graph, i.e., the sequence of candidates that maximizes the total weight utility. Despite its effectiveness, the method struggles with complex road topologies, noisy GPS data, and scalability issues when processing large datasets.
The goal of our work is to investigate how the ST-Matching algorithm can be improved both in terms of its performance efficiency and the quality of its results. To address this, a set of modified versions of the method is proposed, each designed to overcome specific limitations of the original implementation and enhance the overall performance. In particular, the modified ST-Matching is introduced, refining the original method by improving the candidate point selection method, the spatial analysis component, and the temporal component to better reflect realistic movements, travel time, and speed. Moving on to the realm of low-frequency data, we investigate the potential of adding a behavioral score based on historical data to the algorithm, leading to the design of a new behavioral score and the development of the STB-Matching method.
In addition, this paper presents an evaluation framework designed to assess the performance of the modified algorithms. Several metrics are proposed to examine the methods from different perspectives, including computational efficiency, matching accuracy, adherence to topological constraints, and execution speed. Throughout them, a comprehensive comparative analysis is performed to assess the improvements of the new methodologies with respect to the baseline method. The performance and results of thse algorithms are also compared to the baseline on low-frequency data. The proposed approach is validated through a case study in Milan, a key metropolitan area in Northern Italy. The results show clear improvements over the baseline ST-Matching implementation and highlight which modifications are most effective under various operating conditions.
@@ 2.1 Prior Definitions
Considering the the road network graph $G(V,E)$, the path $P_{ij}$ between two vertices $V_{i}$ and $V_{j}$ is a set of connected road segments in $E$, $e_{1}\rightarrow e_{2}\rightarrow\dotsi\rightarrow e_{k}$, where $V_{i}$ is the starting point of the first edge $e_{1}$, and $V_{j}$ is the ending point of the last edge $e_{k}$.
@@ 2.2 Overview of the Baseline Method
The observation probability measures the likelihood that GPS sample point $p_{i}$ corresponds to a candidate point $c_{i}^{s}$, given the Euclidean distance between them. This likelihood assumes that the GPS error has a normal distribution, typically with a zero mean and a standard deviation ($\sigma$), which is set to $\sigma=20$ based on the empirical evaluations [6]. The observation probability is calculated using the probability density function of such a distribution as follows:
@@
After spatial and temporal analysis, for each trajectory $T$, it is possible to create a candidate graph. It is a Trellis graph [11], of which the set of nodes includes the candidate points for each GPS sample $p_{i}$, and the set of edges is made of links connecting candidates of consecutive points. Formally, the candidate graph is defined as $G_{T}=(V_{T},E_{T})$ with:
@@ 2.3 Proposed Modifications to ST-Matching
    Dynamic candidate preparation, which affects the search area for the candidate points using a dynamic buffer and through GPS accuracy. It is discussed in Section 2.3.1;
@@
    Dynamic observation probability, which modifies the first component of the spatial analysis again based on data accuracy, and is presented in Section 2.3.1;
@@
    Redesigned temporal score function, which replaces entirely the temporal function of the original algorithm and is detailed in Section 2.3.2;
@@ 2.3.2 Redesigned Temporal Function
$v_{avg}(i)=\frac{\text{dist}_{e}(p_{i-1},p_{i})}{\Delta t_{obs}(i)},$ (8)
$\Delta t_{est}(i,s,t)=\frac{\text{dist}_{n}(c^{t}_{i-1},c_{i}^{s})}{v_{avg}(i)}.$ (9)
@@
The speed penalty is designed to compare the average speed estimated from the trajectory data $v_{\text{avg}}$, with the average speed limit of the road segments in a candidate path $v_{\text{lim}}$, and penalizes only over-speeding when the estimated average speed goes beyond the average speed limit. The average observed speed $v_{\text{avg}}$ is defined in Equation 8 and the average speed limit $v_{\text{lim}}$ is defined in the following Equation 11 for each candidate pair $c_{i-1}^{t},c_{i}^{s}$ of consecutive points $p_{i-1},p_{i}$:
@@
where $E_{i,s,t}=E_{(i-1,s)\rightarrow(i,t)}$ is the set of edges of the road network included in the shortest path between candidate $c_{i-1}^{t}$ of $p_{i-1}$ and candidate $c_{i}^{s}$ of $p_{i}$. Similarly to the travel time penalty function, the speed penalty is defined with an exponential shape in Equation 12. It keeps the value of 1 when the average speed is below the limit, and drops as it goes beyond the limit.
@@
Final considerations regard the speed variation penalty. Assuming a moving object on the road network of a city, throughout a trajectory, it passes many road segments and intersections, while prefering to follow relatively direct routes rather than changing the road at every intersection. As a result, the speed limit along a real trajectory should not fluctuate dramatically. The speed variation penalty is introduced to capture this behavior. The penalty is defined based on the ratio of the standard deviation to the mean of speed limits along the candidate path. Considering two candidates, $c_{i-1}^{t}$ and $c_{i}^{s}$, and their candidate path $E_{i,s,t}$, we define the mean and standard deviatiton of the speed limits in the candidate path, and thus the speed variation penalty, as follows:
@@
After the calculation of these three components, the new temporal score function is computed as the product of these penalties:
@@ 2.3.3 Behavioral Analysis
The computation of the behavioral score is as follows. Using historical GPS trajectores of a specific period of time, we perform the candidate point selection stage of the algorithm using the dynamic buffer around all the GPS points in the historical dataset. Each candidate point is mapped to a road segment of the network graph. For every instance a candidate point is associated with a given edge, a counter for that edge is incremented by one. As a result, each edge in the network contains a usage score representing how often it has probably been used in the past trajectories. To deal with the wide range and skewness in edge scores, we apply a logarithmic normalization to scale the values between 0 and 1:
@@ 2.3.4 Final Score Functions
As a result of all the analyses and modifications discussed so far, we define the score functions of two versions of the algorithm. The modified ST-Matching uses a score function combining Equations 3 and 16: the buffer and $\sigma$ parameters used to calculate Equation 3 are the dynamic versions, while the temporal function is modified according to Equation 16.
@@ 2.4.1 Efficiency Metrics
The efficiency metrics assess the computational efficiency of the algorithm. By evaluating these metrics, we quantify the overhead introduced by different candidate selection strategies (e.g., fixed vs. dynamic buffers) and identify trade-offs between algorithmic complexity and matching accuracy. Specifically, we look at the following quantities per each trajectory $T$ of the dataset:
@@ 2.4.3 Topology Metrics
  • Number of revisited edges: The number of times the same edge in the road network graph is used more than once for the same trajectory. Frequent revisits are typically physically implausible for standard transit and suggest that the algorithm is “trapped” oscillating between candidate edges.
@@
  • Number of loops: Defined by the number of polygons formed by the matched path. While naturally present in roundabouts, an excess of loops identifies topological instability and physically-implausible tracing caused by high GPS noise or weak temporal constraints.
@@ 2.4.4 Speed-Based Metric
The speed-based metric evaluates how realistic the reconstructed travel speed is compared to the expected speed estimated from the GPS samples. The speed estimated by the GPS data is treated as a reference. The metric is calculated as the speed relative deviation.
@@ 3.1 GPS Trajectory Dataset
The trajectory observations analyzed in this study are located in the urban area of Milan, Italy, and its surroundings and refer to the period from December 1st to December 20th, 2019. The initial dataset is subjected to a filtering process to guarantee the high quality of the location data and its consistency with the objectives of the study. While the focus of the analysis is on the urban area of Milan, the dataset contains many records outside this area, probably related to daily commuters. Therefore, the first filter is to exclude the records outside the urban area, which is done using the official shapefile of the Milan urban area from the open data portal of the municipality [3].
The sampling interval of the GPS records is quite high in most cases, with the average time interval of 16 seconds over the whole dataset, which means that this is a case of high-frequency GPS data. In order to make sure that the trajectories that are potentially going to be used later in Map Matching are long enough, only the ones with at least 10 GPS records throughout the trajectory are kept.
Finally, in this study, the focus is on the vehicular mobility on the drivable road network, which means trajectories corresponding to walking users must preferably be disregarded. This is achieved by calculating an approximate average speed for each trajectory using the direct distance between each two adjacent points and the time interval between them, and keeping the trajectories with an average speed of at least 6km/h to make sure there are no walking users in the dataset. In the end, the final dataset is composed of trajectory IDs corresponding to motor vehicles moving in the city of Milan.
@@ 3.2 Road Network Description
The road network used in this paper covers the urban area of Milan. It is retrieved from OpenStreetMap [8] using the OSMnx Python library [1]. The network type drive is used in this case, which includes only drivable roads. The retrieved graph is directed and multi-edge, meaning multiple edges may exist between a pair of nodes.
@@
The speed attribute of the graph edges showed missing or non-numerical values, a common limitation in the OpenStreetMap database. Since this attribute is essential for the temporal analysis of Map Matching, the missing value imputation is addressed in several steps. First, in cases where there are no numerical values, but instead a road type code, it is possible to infer information about the speed limits for the corresponding code from the OpenStreetMap Wiki [7]. In few cases where there is a list of two values for the speed limit, the greater values are kept. Later, dealing with missing values, assuming that different segments of the same street usually have the same speed limit, we use the name attribute (i.e., the street name) to fill in the missing values if other segments of the street with the same name contain a maximum speed value. Finally, for the rest of the missing values, another attribute called highway (i.e., the highway type) is used considering the mode of maximum speed values for each road type. The final distribution of the speed limits of the network edges comprises a majority of road segments of 30 km/h (56.0%), followrd by many road segments of 50km/h (38.1%), and the remaining with other velocities.
@@ 3.3 Results
To simulate a low-frequency trajectory, a subsample of 100 trajectories is considered and limited: for each trajectory from the first GPS point we keep the poits with at least a 2 minute time interval from the previous one. The low-frequency trajectories are then used to compare ST-Matching with the STB-Matching version including behavioral analysis. The evaluation is performed exploiting the proposed metrics, but also comparing the results obtained with the synthetic low-frequency data with the ones of their original counterpart, which is used as a ground truth. Moreover, for the STB-Matching, the edge score of the road network is computed using a different subsample of size 400 of the trajectory dataset.
@@
This section compares the result of Map Matching between the original ST-Matching and the modified version which includes dynamic buffer, dynamic observation probability, and the new temporal function. The comparison is carried out through the metrics mentioned in Section 2.4. It is based on the distribution of the metric values and the statistical significance expressed by a t-test.
@@
It is observed in Figure 7 that the speed relative deviation, similar to other metrics, signals an improvement in the results and potentially more realistic paths regarding the temporal aspects of the Map Matching process.
@@
The same comparison across all metrics is presented in this section using to evaluate the original method and the STB variant.
First, the edge score of the behavioral analysis is computed with 400 historical trajectories. Figure 10 illustrates the distribution of the edge score on the road network of the urban area of Milan using the available dataset. It can be observed that the roads directed towards the suburban areas and the central rings are generally darker, meaning that they have higher scores and have been used more frequently according to the available dataset.
@@
For a comprehensive comparison, we use the lower-frequency synthetic sample of 100 trajectories of the dataset to assess the main metrics, and their original counterpart as a ground truth. As expected, efficiency metrics are positively affected by the new algorithm due to the dynamic buffer that results in less candidate points and less runtime. The result of the t-test, as written on the graphs in Figure 11, confirms this effect.
@@
The speed relative deviation also does not show any significant differences between the two algorithms. It is clearly observed in Figure 14 that the distribution of the values and the t-test results indicate almost no change in this metric.
@@
Finally, we compare the results of the path enhancement using the synthetic low-frequency data and their original counterpart, assuming the matched path of the latter as the ground truth and checking how close are the results of ST-Matching and STB-Matching using the former. In order to do so, the edges of the final matched paths by ST-Matching and STB-Matching are compared with that of the modified ST-Matching to see the number of edges they have in common. As summarized in Table 3, among 100 trajectories, in 72 cases ST-Matching and STB-Matching had the same number of edges in common with the modified ST-Matching results. STB-Matching demonstrated superior performance in 15% of instances, while ST-Matching proved more effective in 13% of cases. This reveals that improving the performance of ST-Matching algorithm for low frequency GPS data is a demanding job and might need more investigation.
Table 3: Overlap between ST-Matching and STB-Matching using synthetic low-frequency data with the results of the modified ST-Matching on original high-frequency data.